# Learning Robust Precipitation Forecaster by Temporal Frame Interpolation

### 30 Nov 2023 v1

@@ 4.1 Experimental Setting
Weather4Cast’23 Transfer Training Datasets
consist of satellite data and radar ground truth data collected from 7 regions in Europe in 2019 and 2020. The satellite data was obtained from a geostationary meteorological satellite operated by the European Organization for the Exploitation of Meteorological Satellites (EUMETSAT). The radar data is obtained from the Operational Program for Exchange of Weather Radar Information (OPERA) radar network. The satellite data consists of visible, infrared, and water vapor bands modalities (e.g., IR016, IR039, IR087, IR097, IR108, IR120, IR134, VIS006, VIS008, WV062, WV073).
Each spectral band provides different information depending on their absorption and scattering characteristics with atmosphere composition. For instance, VIS channels represent information about albedo, shape, and roughness using solar radiation, but only available during the daytime. IR channels sense re-emitted object radiation and can be converted to brightness temperature, and WV channels represent the amount of water vapor in the mid and upper atmosphere. It is commonly used to detect clouds from the background and analyze cloud properties using single and multi-channel. The satellite image covers a spatial resolution of 12 km x 12 km with 15-minute intervals. The radar image consists of $252\times 252$ pixels, the same as the satellite data, but covers only a central region of the satellite image with a spatial resolution of 2 km x 2 km (target region).
Weather4Cast’23 Transfer Evaluation Datasets
consist of regions (spatial shift) R08, R09, R10 that are not included in the training set, and R15, R34, R76, R04, R05, R06, R07, R08, R09 and R10 in 2021 (temporal shift).
Implementation detail
As is described in section 3.3, we adopt the modified 3D-UNet as our network [1] in the early time of the competition and switch to 2D-Unet in the late time. The best result on the leaderboard is based on 2D-UNet. However, some of the ablations are experimented on 3D-UNet. The batch size of one GPU was set to 8. In addition, 90 epochs were used and 4 positive weights of binary cross entropy were used. The initial learning rate was set to 1e-4, weight decay to 0.02, and dropout rate to 0.4. AdamW was used as the optimizer, and the learning rate was reduced by 0.9 when the loss was higher than the previous epoch validation loss. An early stop criterion is applied if the validation loss All experiments were performed on Nvidia RTX 4090 $\times$ 4 GPUs.
@@ 4.3 Ablation Study
We validate the effectiveness of our methods on the validation set (R15, R34, R76, R4, R5, R6, and R7 in 2019 and 2020) of the core track. Here, list the mean CSI and mean F1 on the core validation set. The components of our methods include Input Cropping (IC) to 126x126, Output Cropping (OC) to 42x42, and geometric Augmentation (Aug), as well as newly proposed Multi-Level Dice loss (ML-Dice) and Temporal Frame Interpolation (TFI).
@@ Results
Table 2 shows the results. We can draw the following conclusions. (1) Impact of Temporal Frame Interpolation (TFI): The top-performing method is UNet2D with a combination of Multi-Level Dice loss (ML-Dice), Input Cropping (IC), Output Cropping (OC), Augmentation (Aug), and Temporal Frame Interpolation (TFI), scoring highest in both mCSI (0.1069) and mF1 (0.1738). The inclusion of TFI in this configuration suggests its significant contribution to improving the model’s performance. (2) Effectiveness of Multi-Level Dice loss (ML-Dice): Models utilizing ML-Dice consistently outperform those using standard Dice loss. This indicates the effectiveness of ML-Dice in improving the model’s accuracy. (3) Comparison of UNet2D and UNet3D Architectures: UNet2D models appear to outperform UNet3D models under similar configurations. This might suggest that the temporal information may not be fully explored by convolution on the temporal dimension. (4) There is a clear trend showing that as more components (IC, OC, Aug) are added, the performance improves, regardless of the UNet version or loss function used. This underlines the importance of these preprocessing and augmentation techniques in enhancing model performance. Each additional component (IC, OC, Aug) contributes to a steady increase in both mCSI and mF1 scores. For example, comparing UNet3D (Dice) with each subsequent addition shows a gradual improvement, from 0.0598 to 0.0932 in mCSI and from 0.0981 to 0.1539 in mF1. (5) The least effective configuration is UNet3D with only Dice loss, while the most effective is UNet2D with the full suite of components including TFI. This wide range in performance across configurations underscores the impact of methodological choices on the model’s predictive capabilities.
@@ 5 Conclusions
Over the past years, accurate predictions of rain events have become ever more critical, with climate change increasing the frequency of unexpected rainfall. Weather4cast 2023 competition aims to build machine learning models with the ability of super-resolution rain movie prediction under spatio-temporal shifts. This year’s metric requires finer-grain prediction of the precipitation, posing an even bigger challenge to the model’s robustness. In this paper, we proposed a solution with a newly proposed temporal frame interpolation that enhances the robustness by creating synthetic between sampling intervals. We also propose the multi-level dice loss that takes into account the ordinal relationship between rainfall rates, improving the prediction accuracy. By combining several explored training strategies and proposed TFI and multi-level dice loss, our model can achieve the best result on the transfer learning leaderboard of Weather4Cast’23 competition, showing the practical applicability and superiority of our method over existing approaches under rigorous and competitive conditions. In conclusion, the Temporal Frame Interpolation method, supported by a carefully designed loss function and data augmentation strategy, offers a promising direction for future developments in high-dimensional weather forecasting tasks. We hope that our contributions will serve as a stepping stone for subsequent innovations in the field.

### 01 Dec 2023 v2

@@ 4.1 Experimental Setting
The datasets used in Weather4cast’23 do not change compared to Weather4cast’22.
Training set.
The training set encompasses both satellite and terrestrial radar data, encompassing a span of two years, 2019 and 2020, across seven distinct European regions. The satellite data was sourced from a geostationary meteorological satellite managed by EUMETSAT, while ground truth radar data was acquired from OPERA, a collaborative weather radar information exchange program. This satellite dataset includes a variety of spectral bands, capturing information across the visible spectrum, infrared, and water vapor channels, such as IR016 and VIS006, among others. Each spectral frequency provides unique insights due to their distinct interactions with the atmospheric constituents. For example, the visible (VIS) bands harness solar radiation to yield data on albedo and surface texture but are limited to daylight hours for data collection. In contrast, infrared (IR) bands detect radiation re-emitted from objects, which can be translated into measures of brightness temperature, while water vapor (WV) bands quantify the water vapor present in the higher atmospheric layers. These bands are instrumental for differentiating clouds from the surroundings and examining cloud dynamics via single or combined channel analysis. Satellite imagery has a spatial footprint of 12 km x 12 km, captured at quarter-hourly intervals, whereas radar imagery, with a resolution of 2 km x 2 km, focuses on a central subset of the satellite’s coverage.
Transfer test set.
The test set for the transfer leaderboard consists of regions not present in the training dataset (indicating a spatial shift), labeled R08, R09, and R10, in addition to a set of regions from the year 2021 (indicating a temporal shift), including R15, R34, R76, R04, R05, R06, R07, R08, R09, and R10.
Core validation set.
The core validation set consists of data from all the seen regions (R15, R34, R76, R4, R5, R6, and R7) in 2019 and 2020.
Details on Implementation
As referenced in section 3.3, our early competition strategy involved utilizing a variant of the 3D-UNet architecture [1], which later transitioned to a 2D-UNet model. It was this latter model that yielded the leading results on the leaderboard. However, certain ablation studies were conducted using the 3D-UNet configuration. We configured our single GPU batch size at eight and ran the models for a total of 90 epochs. The initial learning rate was established at 0.0001, with a weight decay set at 0.02 and a dropout rate maintained at 0.4. The optimization was managed via AdamW, featuring a learning rate scaled down by a factor of 0.9 whenever the validation loss exceeded that of the previous epoch. Furthermore, an early stopping mechanism was enforced to halt training if the validation loss plateaued. All experiments were executed on 4 Nvidia RTX 4090 GPUs.
@@ 4.3 Ablation Study
We validate the effectiveness of our methods on the validation set (R15, R34, R76, R4, R5, R6, and R7 in 2019 and 2020) of the core track. Here lists the mean CSI and mean F1 on the core validation set. The components of our methods include Input Cropping (IC) to 126x126, Output Cropping (OC) to 42x42, and geometric Augmentation (Aug), as well as newly proposed Multi-Level Dice loss (ML-Dice) and Temporal Frame Interpolation (TFI).
@@ Results
Table 2 shows the results. We can draw the following conclusions. (1) Impact of Temporal Frame Interpolation (TFI): The top-performing method is UNet2D with a combination of Multi-Level Dice loss (ML-Dice), Input Cropping (IC), Output Cropping (OC), Augmentation (Aug), and Temporal Frame Interpolation (TFI), scoring highest in both mCSI (0.1069) and mF1 (0.1738). The inclusion of TFI in this configuration suggests its significant contribution to improving the model’s performance. (2) Effectiveness of Multi-Level Dice loss (ML-Dice): Models utilizing ML-Dice consistently outperform those using standard Dice loss. This indicates the effectiveness of ML-Dice in improving the model’s accuracy. (3) Comparison of UNet2D and UNet3D Architectures: UNet2D models appear to outperform UNet3D models under similar configurations. This might suggest that convolution on the temporal dimension may not fully explore the temporal information. (4) A clear trend shows that as more components (IC, OC, Aug) are added, the performance improves, regardless of the UNet version or loss function used. This underlines the importance of these preprocessing and augmentation techniques in enhancing model performance. Each additional component (IC, OC, Aug) contributes to a steady increase in both mCSI and mF1 scores. For example, comparing UNet3D (Dice) with each subsequent addition shows a gradual improvement, from 0.0598 to 0.0932 in mCSI and from 0.0981 to 0.1539 in mF1. (5) The least effective configuration is UNet3D with only Dice loss, while the most effective is UNet2D with the full suite of components including TFI. This wide range in performance across configurations underscores the impact of methodological choices on the model’s predictive capabilities.
@@ 5 Conclusions
Over the past years, accurate predictions of rain events have become ever more critical, with climate change increasing the frequency of unexpected rainfall. Weather4cast 2023 competition aims to build machine learning models with the ability of super-resolution rain movie prediction under spatio-temporal shifts. This year’s metric requires finer-grain prediction of the precipitation, posing an even bigger challenge to the model’s robustness. In this paper, we propose a solution with a newly proposed temporal frame interpolation that enhances the robustness by creating synthetic between sampling intervals. We also propose the multi-level dice loss that takes into account the ordinal relationship between rainfall rates, improving the prediction accuracy. By combining several explored training strategies and proposed TFI and multi-level dice loss, our model can achieve the best result on the transfer learning leaderboard of Weather4cast’23 competition, showing the practical applicability and superiority of our method over existing approaches under rigorous and competitive conditions. In conclusion, the Temporal Frame Interpolation method, supported by newly proposed ML-Dice loss and data augmentation strategy, offers a promising direction for future developments in high-dimensional weather forecasting tasks. We hope that our contributions will serve as a stepping stone for subsequent innovations in the field.